OpenStudy (anonymous):

use l'hopital's rule to find the limit x gos to inf xsin(19/x)

OpenStudy (anonymous):

Maybe read the question more carefully. \[\lim_{x \rightarrow \infty} x \sin (19/x)\]doesn't look like a case where you'd use L'Hopital's rule.

OpenStudy (amistre64):

sin / (1/x) perhaps

OpenStudy (anonymous):

can you explain

OpenStudy (amistre64):

or ... x / (csc(19/x))

OpenStudy (amistre64):

you are trying to resturcture the setup to a 0/0 form

OpenStudy (anonymous):

@amistre64 I like your first one better

OpenStudy (amistre64):

the first one might be good since we got a 19/x :) just wasnt all that sure

OpenStudy (anonymous):

were did the 19 go in the first one

OpenStudy (anonymous):

This is what she meant: \[\lim_{x \rightarrow \infty} \sin(19/x)/(1/x)\] Take the derivative of the top and the derivative of the bottom. Some (1/x^2)s cancel. limit as cos(19/x) goes to zero is one The answer should be 19.

OpenStudy (amistre64):

\[x~sin(kx)\] \[\frac{sin(kx)}{1/x}\] \[\frac{\cancel{cos(k/x)}^1*~\bcancel{-}k\cancel{/x^2}}{\cancel{-1/x^2}}=\]

OpenStudy (anonymous):

@amistre64 :)

OpenStudy (anonymous):

thanks